how to solve the function of demand and supply good are
QD=50-2P
QS=4P+2P
QD=QS
1
Expert's answer
2020-11-03T11:18:58-0500

"Q_{D}"=50-2P--------(1)

"Q_{S}" = 4P+ 2P--------(2)-

"Q_{D}=Q_{S}" --------------(3)

Put the value of QD and Qs in eqn (3)

50-2P=4P+2P

8P = 50

P = 6.25

where P is the equilibrium Price.

Put the value of P in eqn (1) and (2),

we get,

"Q_{D}" ="50-2\times P" ="50-2\times6.25=37.5"

"Q_{S}" ="6\times P" ="6\times6.25=37.5"

At point P where the function of demand goods are equal to the function of supply goods.